# Generalized Nash Equilibrium Solutions in Dynamic Games With Shared Constraints

### v2

@@ 5 Analytical Examples
The KKT conditions of the problem are:
$\begin{split}&\nabla_{x}\mathcal{L}_{1}=0\Rightarrow 2x-2+\alpha_{1}\sigma=0\Rightarrow x=1-\frac{1}{2}\alpha_{1}\sigma\\ &\nabla_{y}\mathcal{L}_{2}=0\Rightarrow 2y-1+\alpha_{2}\sigma=0\Rightarrow y=\frac{1}{2}-\frac{1}{2}\alpha_{2}\sigma\\ &0\leq\sigma\perp(x+y-1)\leq 0\\ &\sigma\geq 0,(x+y-1)\leq 0\end{split}$ (22)
There now two options according to the complementary slackness condition. The first option is that the constraint is inactive. This option gives a solution of $x=1,y=\frac{1}{2}$ that violates the constraints and is therefore eliminated. The second option is that the constraint is active:
$\begin{split}&x+y=1\Rightarrow 1-\frac{1}{2}\alpha_{1}\sigma+\frac{1}{2}-\frac{1}{2}\alpha_{2}\sigma=1\\ &\Rightarrow\sigma=\frac{1}{\alpha_{1}+\alpha_{2}}\\ \end{split}$ (23)
According to Theorem 2 for every given $\alpha_{1}$ and $\alpha_{2}$ $x^{*}$ and are NE:
@@
$\alpha=\frac{\alpha_{1}}{\alpha_{1}+\alpha_{2}}\Rightarrow 1-\alpha=\frac{\alpha_{2}}{\alpha_{1}+\alpha_{2}}\Rightarrow\left\{\begin{aligned} &x^{*}=1-\frac{1}{2}\alpha\\ &y^{*}=\frac{\alpha}{2}\\ \end{aligned}\right.$ (25)
Since $\alpha_{1}>0$ and $\alpha_{2}>0$ it suggests that $\alpha\in(0,1)$. The solution for $\alpha\rightarrow 0$ gives the solution that converges to $(x^{*}=1,y^{*}=0)$ which is the unconstrained solution for player 1 and player 2 just responds - Stackelberg Equilibrium where player 1 is the leader and player 2 is the follower. The solution for $\alpha\rightarrow 1$ gives the solution the converges to $(x^{*}=\frac{1}{2},y^{*}=\frac{1}{2})$ which is the unconstrained solution for player 2 and player 1 just responds - Stackelberg Equilibrium where player 2 is the leader and player 1 is the follower. For $\alpha=\frac{1}{2}$ the solution corresponds to the normalized solution $(x^{*}=\frac{3}{4},y^{*}=\frac{1}{4})$.
The second example given here is a dynamic game presented in 2.1. Applying the technique presented in this paper the and notating the factor as $\alpha_{1},\alpha_{2},\alpha_{3}$ for the shared constraint, the Lagrangians are:
$\begin{split}\mathcal{L}_{1}=-x_{1}+x_{2}+\frac{1}{2}v_{1}^{2}+\mu_{1}(x_{1}-x_{1}(0)-\Delta t\cdot v_{1})+\alpha_{1}\sigma(x_{2}-x_{3})\\ \mathcal{L}_{2}=-x_{2}+x_{1}+\frac{1}{2}v_{2}^{2}+\mu_{1}(x_{2}-x_{2}(0)-\Delta t\cdot v_{2})+\alpha_{2}\sigma(x_{2}-x_{3})\\ \mathcal{L}_{3}=-x_{1}+x_{2}+\frac{1}{2}v_{3}^{2}+\mu_{1}(x_{3}-x_{3}(0)-\Delta t\cdot v_{3})+\alpha_{3}\sigma(x_{2}-x_{3})\end{split}$ (26)
Applying the stationary condition:
$\begin{split}&\frac{\partial\mathcal{L}_{1}}{\partial[x_{1},v_{1}]}=\begin{bmatrix}-1+\mu_{1}\\ v_{1}-\mu_{1}\Delta t\end{bmatrix}=0\Rightarrow\mu_{1}=1\Rightarrow v_{1}=\Delta t\\ &\frac{\partial\mathcal{L}_{2}}{\partial[x_{2},v_{2}]}=\begin{bmatrix}-1+\mu_{2}+\alpha_{2}\sigma\\ v_{2}-\mu_{2}\Delta t\end{bmatrix}=0\Rightarrow\mu_{2}=1-\alpha_{2}\sigma\Rightarrow v_{2}=\Delta t(1-\alpha_{2}\sigma)\\ &\frac{\partial\mathcal{L}_{3}}{\partial[x_{3},v_{3}]}=\begin{bmatrix}\mu_{3}-\alpha_{3}\sigma\\ v_{3}-\mu_{3}\Delta t\end{bmatrix}=0\Rightarrow\mu_{3}=\alpha_{3}\sigma\Rightarrow v_{3}=\Delta t\alpha_{3}\sigma\end{split}$ (27)
By the complementary slackness condition, there are two options for a solution, either the constraint is active ($\sigma=0$), which gives the following solution:
$(x_{1},x_{2},x_{3})=(x_{1}(0)+\Delta t^{2},x_{2}(0)+\Delta t^{2},x_{3}(0))=(1,1.5,1)$ (28)
This solution violates the shared constraint and is therefore eliminated. The second option is that the shared constraint is active $(\sigma>0)$:
$\begin{split}&x_{2}=x_{3}\Rightarrow x_{2}(0)+\Delta t\cdot v_{2}=x_{3}(0)+\Delta t\cdot v_{3}\Rightarrow\\ &x_{2}(0)+\Delta t^{2}(1-\alpha_{2}\sigma)=x_{3}(0)+\Delta t^{2}\alpha_{3}\sigma\Rightarrow\sigma=\frac{\Delta t^{2}-x_{3}(0)+x_{2}(0)}{(\alpha_{2}+\alpha_{3})\Delta t^{2}}\Rightarrow\\ &(x_{1},x_{2},x_{3})=(x_{1}(0)+\Delta t^{2},\frac{\alpha_{3}x_{2}(0)+\alpha_{2}x_{3}(0)+\alpha_{3}\Delta t^{2}}{\alpha_{2}+\alpha_{3}},\frac{\alpha_{3}x_{2}(0)+\alpha_{2}x_{3}(0)+\alpha_{3}\Delta t^{2}}{\alpha_{2}+\alpha_{3}}\end{split}$ (29)
@@
$\alpha=\frac{\alpha_{2}}{\alpha_{2}+\alpha_{3}}\Rightarrow 1-\alpha=\frac{\alpha_{3}}{\alpha_{2}+\alpha_{3}}\Rightarrow\left\{\begin{aligned} &x^{*}_{1}=1\\ &x^{*}_{2}=\alpha x_{3}(0)+(1-\alpha)(\Delta t^{2}+x_{2}(0))=\frac{3}{2}-\frac{3}{4}\alpha\\ &x^{*}_{3}=\alpha x_{3}(0)+(1-\alpha)(\Delta t^{2}+x_{2}(0))=\frac{3}{2}-\frac{3}{4}\alpha\end{aligned}\right.$ (30)
@@
Since $\alpha_{2}>0$ and $\alpha_{3}>0$, it suggests that $\alpha\in(0,1)$. This gives all the solutions for $x_{2}=x_{3}\in(0.75,1.5)$.
6 Selecting a GNE Point Among Many
@@
In previous section we have shown how to utilize an MCP formulation to calculate a set of possible GNE of a GNEP. To choose a solution from the possible GNE set, a bi-level optimization scheme is suggested in this paper. The optimization problem we proposes chooses the set of factor matrices $\{A_{i}\}_{i=1}^{N}$ which that minimize a given cost function:
$\begin{split}&\min_{A_{i},..,A_{N}\in\mathcal{D}^{+}_{m}}{J_{0}(x^{*})}\\ &s.t.\\ &(x^{i})^{*}\in\mathcal{S}_{i}(x^{-i},A_{1},...,A_{N})\quad\forall i\in\{1,...,N\}\end{split}$ (31)
@@
In general, the bi-level optimization 31 can be hard to solve. The cost to minimize can be of any form and can try to achieve different goals. A few possible objectives include: 1) minimize the sum of all the players’ costs. 2) Optimize for the cost of one of the players - this will lead to a strategy close to a Stackelberg Equilibrium (in a 2 players scenario). 3) Optimize over some property of the game - maximize the average velocity in a racing game or maximize the interaction between players. To see the effect of such bi-level optimization, the problems presented in section 5 are solved using cost 1):
$J_{0}=\sum_{i=1}^{N}{J_{i}(x^{i},x^{-i})}$ (32)
@@
In the case of the first example, the bi-level optimization is:
$\begin{split}\min_{0<\alpha<1}&{(x^{*}-1)^{2}+(y^{*}-\frac{1}{2})^{2}}\\ &s.t.\\ &x^{*}=1-\frac{1}{2}\alpha\\ &y^{*}=\frac{\alpha}{2}\end{split}$ (33)
@@
$\begin{split}&J_{0}(\alpha)=(1-\frac{1}{2}\alpha-1)^{2}+(\frac{\alpha}{2}-\frac{1}{2})^{2}\\ &\Rightarrow J_{0}(\alpha)=\frac{1}{4}(2\alpha^{2}-\alpha+1)\\ &\Rightarrow\frac{dJ_{0}}{d\alpha}=\alpha-1\end{split}$ (34)
By setting $\frac{dJ_{0}}{d\alpha}=0$, the optimal solution given the chosen $J_{0}$ is $\alpha^{*}=1$. Even though the domain of $\alpha\in(0,1)$ technically doesn’t include $1$, in this case $\alpha=1$ gives a feasible solution:
$x=y=\frac{1}{2}$ (35)
Recall that the normalized solution approach gives a different solution $(x=\frac{3}{4},y=\frac{1}{4})$. For the same choise of $J_{0}$ type, the second example give the following:
$\begin{split}\min_{0<\alpha<1}&{-x^{*}_{1}+x^{*}_{2}+\frac{(v_{1}^{*})^{2}+(v_{2}^{*})^{2}+(v_{3}^{*})^{2}}{2}}\\ &s.t.\\ &x^{*}_{1}=1\\ &x^{*}_{2}=(1-\alpha)(\Delta t^{2}+x_{2}(0))+\alpha x_{3}(0)+(1-2\alpha)\Delta t^{2}\\ &v_{1}^{*}=\Delta t\\ &v_{2}^{*}=\frac{(1-\alpha)\Delta t^{2}+\alpha(x_{3}(0)-x_{2}(0))}{\Delta t}\\ &v_{3}^{*}=\frac{(1-\alpha)\Delta t^{2}-(1-\alpha)(x_{3}(0)-x_{2}(0))}{\Delta t}\end{split}$ (36)
Substituting the terms for $(x_{1},v_{1},x_{2},v_{2})^{*}$ into $J_{0}$ and taking the derivative with respect to $\alpha$:
$\frac{dJ(\alpha)}{d\alpha}=0$ (37)
where, Solving for the optimal $\alpha^{*}$ gives a value greater than the maximal allowed, so the optimal value is $\alpha^{*}=1$ and this corresponds to the solution:
@@
7 Numerical Two-Car Racing Problem
@@
One of the most popular ways to solve a problem of autonomous car racing is by using a DG approach. In a 2 cars race, each car has a unique dynamic model and a cost function to minimize. A DG formulation is set up for each car with some shared prediction horizon. The shared constraint of the game is Collision Avoidance (CA) between both players. The DG is solved by finding the GNE of the problem, usually using a MCP formulation solved with a numerical solver, and the shared constraints Lagrange multipliers are assumed to be the same [2, 8]. As in the problem presented in section 2.1, this problem usually has more than one GNE and there is no good reason to prefer the normalized solution. The kinematic bicycle model for each car, with the following states:
$x^{i}_{k}=[v,\psi,s,t,X,Y],\quad i=1,2\quad k=0,...,N$ (39)
@@
where, $v$ is the total velocity, $\psi$ is the angular state, and $s,v_{t}$ are the arc-length position of the car in the local Frenet reference frame relative to the center line of the track and lateral deviation from the center-line, respectively. $X$ and $Y$ are the inertial position of the car. $N$ is the prediction horizon. The inputs are:
$u^{i}_{k}=[u_{a},u_{\delta}],\quad i=1,2\quad k=0,...,N-1$ (40)
@@
where, $u_{a}$ is the longitudinal acceleration and $u_{\delta}$ is the steering of the front wheels. To illustrate the effectiveness of the proposed method, the same dynamic model and cost function is used for both players. The Dynamic Game formulation is given by:
$i\in\{1,2\}\left\{\begin{aligned} &\min_{x^{i},u^{i}}-s^{i}_{N}+s^{-i}_{N}+\frac{\beta}{2}\sum_{k=0}^{N-1}\|u^{i}_{k}\|^{2}\\ &\text{s.t.}\\ &x^{i}_{k+1}=f(x^{i}_{k},u^{i}_{k}),\quad k=0,\dots,N-1\\ &x^{i}_{k}\in\mathcal{X},\quad u^{i}_{k}\in\mathcal{U},\quad k=0,\dots,N-1\\ &(X^{1}_{k}-X^{2}_{k})^{2}+(Y^{1}_{k}-Y^{2}_{k})^{2}\geq d_{safe},\quad k=0,\dots,N-1\end{aligned}\right.$
@@
where, $f$ is the dynamics function, $\mathcal{X}$ is the feasible set of the states, $\mathcal{U}$ is the feasible set of the inputs, and $d_{safe}$ is the minimal distance between both vehicles. $\beta=10^{-1}$. As can be seen, both cars have the same properties and the same cost function properties. Each car aims to maximize its own progress while minimizing control effort. A detailed description of the dynamics function $f$ and the feasible sets of states and inputs can be found in the Appendix 9.
@@
To find the GNE of this DG, an MCP formulation 11 is used. To define the Lagrangian of each car by the method presented in 4, a different factor should be given to every CA constraint. In this problem, there are $N$ Constraints for each car. To make the solution more trackable, the number of parameters of the optimization problem is lowered. The factor matrix of the first car is chosen to be $I_{m_{0}}$ as explained in section 4 and the factor matrix of car 2 is reduced to $\alpha I_{m_{0}}$ which can further reduced to a scalar $\alpha$. This means that there is a single parameter that factors the shared constraints Lagrange multipliers:
$L_{1}=-s^{1}_{N}+s^{2}_{N}+\frac{\beta}{2}\sum_{k=0}^{N-1}{\|u^{1}_{k}\|^{2}}+\mu_{1}^{T}\cdot h_{1}(x^{1},u^{1})+\lambda_{1}^{T}\cdot g_{1}(x^{1},u^{1})+\sigma^{T}\cdot g_{CA}(x^{1},x^{2})$
$L_{2}=-s^{2}_{N}+s^{1}_{N}+\frac{\beta}{2}\sum_{k=0}^{N-1}{\|u^{2}_{k}\|^{2}}+\mu_{2}^{T}\cdot H_{2}(x^{2},u^{2})+\lambda_{2}^{T}\cdot g_{2}(x^{2},u^{2})+\alpha\sigma^{T}\cdot g_{CA}(x^{1},x^{2})$
@@
where $\mu_{1},\mu_{2}$ are the Lagrange multipliers associated with the dynamics (equality) constraints $h_{1},h_{2}$. $\lambda_{1},\lambda_{2}$ are the Lagrange multipliers associated with the state and input constraints $g_{1},g_{2}$. $\sigma$ is the Lagrange multiplier vector associated with the CA constraint $g_{CA}$. The next step to solve the MCP formulation is to apply the stationary condition on each Lagrangian and complementary slackness conditions:
$\frac{\partial L_{1}}{\partial(x^{1},u^{1})}=0,\quad\frac{\partial L_{2}}{\partial(x^{2},u^{2})}=0$

### v3

@@ V-A Example 1
Applying the KKT conditions and according to Theorem 2 for every given $\alpha_{1}$ and $\alpha_{2}$, the solutions for $x^{*}$ and $y^{*}$ are NE:
@@
[…]
Since $\alpha_{1}>0$ and $\alpha_{2}>0$ it suggests that $\alpha\in(0,1)$. The solution for $\alpha\rightarrow 0$ gives the solution that converges to $(x^{*}=1,y^{*}=0)$ which is the unconstrained solution for player 1 and player 2 just responds - Stackelberg Equilibrium where player 1 is the leader and player 2 is the follower. The solution for $\alpha\rightarrow 1$ gives the solution the converges to $(x^{*}=\frac{1}{2},y^{*}=\frac{1}{2})$ which is the unconstrained solution for player 2 and player 1 just responds - Stackelberg Equilibrium where player 2 is the leader and player 1 is the follower. For $\alpha=\frac{1}{2}$ the solution corresponds to the normalized solution $(x^{*}=\frac{3}{4},y^{*}=\frac{1}{4})$. The proposed method allows to compute all possible GNEs, whose existence was mentioned in [facchinei2007generalized, facchinei2009generalized]. The second example given here is a dynamic game presented in (II-A). Applying the technique presented in this paper, notating the factor as $\alpha_{1},\alpha_{2},\alpha_{3}$ for the shared constraints factors, and applying the KKT condition gives the following solution:
$\begin{split}x_{1}&=x_{1}(0)+\Delta t^{2}\\ x_{2}=x_{3}=&\frac{\alpha_{3}x_{2}(0)+\alpha_{2}x_{3}(0)+\alpha_{3}\Delta t^{2}}{\alpha_{2}+\alpha_{3}}\end{split}$ (30)
@@
[…]
@@
Since $\alpha_{2}>0$ and $\alpha_{3}>0$, it suggests that $\alpha\in(0,1)$. This gives all the solutions for $x_{2}=x_{3}\in(0.75,1.125)$. From the equilibrium conditions, it follows that these are all GNEs of the given GNEP.
V-B Example 2
Another example presented here is Harker’s problem solved in [harker1991generalized, nabetani2011parametrized]: We consider the following two-player Generalized Nash Equilibrium Problem (GNEP):
Player 1: $\min_{x_{1}}\;x_{1}^{2}+\frac{8}{3}x_{1}x_{2}-34x_{1}$
$\text{s.t.}\;0\leq x_{1}\leq 10$
$x_{1}+x_{2}\leq 15,$
Player 2: $\min_{x_{2}}\;x_{2}^{2}+\tfrac{5}{4}x_{1}x_{2}-24.25x_{2}$
$\text{s.t.}\;0\leq x_{2}\leq 10$
$x_{1}+x_{2}\leq 15.$
Our proposed solution procedure involves defining the Lagrangian for each player, introducing a scaling parameter for the shared constraint:
[…]
where $\sigma$ is the unscaled shared Lagrange multiplier, and $\alpha_{1}$, $\alpha_{2}$ are the scaling factors for players 1 and 2, respectively. Without loss of generality, we set $\alpha_{1}=1$, yielding:
[…]
where $\alpha=\alpha_{2}\in(0,\infty]$. The KKT conditions for the problem are:
$\frac{\partial L_{1}}{\partial x_{1}}=0,\quad\frac{\partial L_{2}}{\partial x_{2}}=0,$ (34)
[…]
$\lambda_{1}^{1},\lambda_{1}^{2},\lambda_{2}^{1},\lambda_{2}^{2},\sigma\geq 0.$
The stationary conditions are:
[…]
One possible solution arises when all constraints are inactive (i.e., all Lagrange multipliers are zero):
$\begin{aligned} &2x_{1}+\frac{8}{3}x_{2}-34=0,\\ &2x_{2}+\frac{5}{4}x_{1}-24.25=0,\end{aligned}\quad\Rightarrow\quad x_{1}^{*}=5,\quad x_{2}^{*}=9.$ (36)
This yields one valid solution. Other solutions arise when at least one constraint is active. For example, if the shared constraint is active ($\sigma>0$), then:
$x_{1}+x_{2}=15.$ (37)
Substituting into the stationarity conditions:
[…]
Solving this system yields:
$\sigma^{*}(\alpha)$ $=\frac{8}{8\alpha-9},$ (39)
$x_{1}^{*}(\alpha)$ $=\frac{72\alpha-69}{8\alpha-9},$
$x_{2}^{*}(\alpha)$ $=\frac{48\alpha-66}{8\alpha-9}.$
Since $\sigma>0$, we require $\alpha\in\left(\frac{9}{8},\infty\right]$. Imposing the constraint $x_{1}\leq 10$ further restricts $\alpha\in\left(\frac{21}{8},\infty\right]$.
This example demonstrates the analytical tractability and flexibility of the proposed method. By introducing scaling parameters on shared constraints, we can derive closed-form expressions for the entire set of GNE solutions, even in nonlinear and non-convex settings. Compared to the parametrized VI approach, our method offers clearer guarantees of solution validity, reduced dimensionality, and broader compatibility with existing solvers. We believe this framework opens new avenues for efficiently solving complex multi-agent optimization problems.
VI Selecting a GNE Point Among Many
@@
In previous section we have shown how to utilize the KKT condition either by directly solving them or by using an MCP formulation to calculate a set of possible GNE of a GNEP. To choose a solution from the possible GNE set, a bi-level optimization scheme is suggested in this paper. The optimization problem we propose chooses the set of factor matrices $\{A_{i}\}_{i=1}^{N}$ which minimize a given cost function:
$\begin{split}&\min_{A_{i},..,A_{N}\in\mathcal{D}^{+}_{m}}{J_{0}(x^{*})}\\ &s.t.\quad(x^{i})^{*}\in\mathcal{S}_{i}(x^{-i},A_{1},...,A_{M})\quad\forall i\in\{1,...,M\}\end{split}$ (40)
@@
In general, the bi-level optimization (40) can be hard to solve. The cost to minimize can be of any form and can try to achieve different goals. A few possible objectives include: 1) minimize the sum of all the players’ costs. 2) Optimize for the cost of one of the players - this will lead to a strategy close to a Stackelberg Equilibrium (in a 2 players scenario). 3) Optimize over some property of the game, for instance maximize the average velocity in a racing game or maximize the interaction between players (to have a game which is fun to watch). To see the effect of such bi-level optimization, the problems presented in section V are solved by minimizing the sum of all the players’ costs:
$J_{0}=\sum_{i=1}^{M}{J_{i}(x^{i},x^{-i})}$ (41)
@@
In the case of the first example, the bi-level optimization becomes a one-layer optimization since the GNE solution of the DG is known (29):
$\begin{split}\min_{0\leq\alpha\leq 1}&{(x^{*}-1)^{2}+(y^{*}-\frac{1}{2})^{2}}\\ s.t.~&x^{*}=1-\frac{1}{2}\alpha,\quad y^{*}=\frac{\alpha}{2}\end{split}$ (42)
@@
$\begin{split}&J_{0}(\alpha)=\frac{1}{4}(2\alpha^{2}-\alpha+1)\Rightarrow\frac{dJ_{0}}{d\alpha}=\alpha-1\end{split}$ (43)
By setting $\frac{dJ_{0}}{d\alpha}=0$, the optimal solution given the chosen $J_{0}$ is $\alpha^{*}=1$. Note that, in general, one should exclude boundary solution of $\alpha$ (i.e. 0 and $1$) as joint constraints might not be satisfied by one player. In this example $\alpha=1$ yields $x=y=\frac{1}{2}$ which is a feasible solution. Recall that the normalized solution approach gives a different solution $(x=\frac{3}{4},y=\frac{1}{4})$. For the same type of high level cost function, the example presented in (II-A) results in the following optimization:
[…]
Substituting the terms for $(x_{1}^{*},v_{1}^{*},x_{2}^{*},v_{2}^{*})$ into $J_{0}$ and solving for the optimal $\alpha^{*}$ gives $\alpha^{*}=1$ and this corresponds to the solution:
@@
VII Two-Cars Racing Problem
@@
One of the most popular approaches to solving an autonomous car racing problem is by using a Dynamic Game (DG) approach. In a 2 cars race, each car has a unique dynamic model and a cost function to minimize. A DG formulation is set up for each car with some shared prediction horizon. The game’s shared constraint is Collision Avoidance (CA) between both players. The DG is solved by finding the GNE of the problem, usually using a MCP formulation solved with a numerical solver, and the shared constraints Lagrange multipliers are assumed to be the same [zhu2023sequential, liu2023learning]. This problem often has multiple GNEs, and there is no clear reason to prefer the normalized solution over others.
A kinematic bicycle model for each car is used, with the following states and inputs:
$x^{i}_{k}=[v,\psi,s,t,X,Y],\quad i=1,2\quad k=0,.,N$ (46)
$u^{i}_{k}=[u_{a},u_{\delta}],\quad i=1,2\quad k=0,.,N-1$
@@
where, $v$ represents the total velocity, $\psi$ is the angular state relative to the center-line, and $s,v_{t}$ are the arc-length position of the car in the local Frenet reference frame relative to the center line of the track and lateral deviation from the center-line, respectively. $X$ and $Y$ are the inertial position of the car. $u_{a}$ is the longitudinal acceleration and $u_{\delta}$ is the steering of the front wheels $N=10$ is the prediction horizon. Finally, the time step is set to $\Delta t=0.1[sec]$.
@@
To illustrate the effectiveness of the proposed method, the same dynamic model and cost function is used for both players. The Dynamic Game formulation for player $i\in\{1,2\}$ is given by:
$\min_{x^{i},u^{i}}-s^{i}_{N}+s^{-i}_{N}+\frac{\beta}{2}\sum_{k=0}^{N-1}\|u^{i}_{k}\|^{2}$
$\text{s.t.}\quad x^{i}_{k+1}=f(x^{i}_{k},u^{i}_{k}),\quad x^{i}_{k}\in\mathcal{X},\quad u^{i}_{k}\in\mathcal{U},$
$(X^{1}_{k}-X^{2}_{k})^{2}+(Y^{1}_{k}-Y^{2}_{k})^{2}\geq d_{safe},\quad k=1,\dots,N-1,$
@@
where, $f$ is the dynamics function, $\mathcal{X}$ is the feasible set of the states, $\mathcal{U}$ is the feasible set of the inputs, and $d_{safe}=0.4[m]$ is the minimal distance between both vehicles, and $\beta=10^{-1}$. As can be seen, both cars have the same characteristics and the same cost function properties. Each car aims to maximize its relative progress while minimizing control effort.
@@
To compute the GNE of this DG, an MCP formulation is numerically solved using the PATH solver [dirkse1995path]. To define the Lagrangian of each car using the method presented in section IV, a different factor should be given to each collision avoidance (CA) constraint. In this problem, there are $N$ shared constraints for each car. To make the solution more tractable, the number of factors is lowered. The factor matrix of the first car is chosen to be $I_{N}$ as explained in section IV, and the factor matrix of car 2 is reduced to $\alpha I_{N}$ which can further reduced to a scalar $\alpha$. The physical interpretation of $\alpha\in\mathbb{R}_{++}$ is the relative aggressiveness of player 2 relative to player 1. If $\alpha<1$ player 2 is more aggressive than player 1, and for $\alpha>1$ player 1 is more aggressive than player 2.
This means that there is a single parameter that factors the shared constraints Lagrange multipliers:
$\mathcal{L}_{1}$ $=-s^{1}_{N}+s^{2}_{N}+\frac{\beta}{2}\sum_{k=0}^{N-1}{\|u^{1}_{k}\|^{2}}+\mu_{1}^{T}\cdot h(x^{1},u^{1})+$ (47)
$\qquad\qquad+\lambda_{1}^{T}\cdot g(x^{1},u^{1})+\sigma^{T}\cdot g_{CA}(x^{1},x^{2})$
$\mathcal{L}_{2}$ $=-s^{2}_{N}+s^{1}_{N}+\frac{\beta}{2}\sum_{k=0}^{N-1}{\|u^{2}_{k}\|^{2}}+\mu_{2}^{T}\cdot h(x^{2},u^{2})+$
$\qquad\qquad+\lambda_{2}^{T}\cdot g(x^{2},u^{2})+\alpha\cdot\sigma^{T}\cdot g_{CA}(x^{1},x^{2})$
@@
where $\mu_{1},\mu_{2}$ are the Lagrange multipliers associated with the dynamics (equality) constraints $h$. $\lambda_{1},\lambda_{2}$ are the Lagrange multipliers associated with the state and input constraints $g$. $\sigma$ is the fictitious Lagrange multiplier vector associated with the CA constraint $g_{CA}$. The next step is to apply the stationary condition on each Lagrangian and complementary slackness conditions:
$\frac{\partial\mathcal{L}_{1}}{\partial(x^{1},u^{1})}=0,\quad\frac{\partial\mathcal{L}_{2}}{\partial(x^{2},u^{2})}=0$ (48)
@@
Solving the above KKT equations for any value of $\alpha\in(0,\infty)$ gives a valid GNE solution.
VII-A Quantitative Analysis of Tuning Aggressiveness
The solution set of two examples are presented here to illustrate the effectiveness and added value of our method. Each example corresponds to a specific initial state of the cars on different tracks.